Circular No. 5897

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993af, 1993ag, 1993ah The European Southern Observatory/European Union student visi-

tors’ group to the La Silla Observatory (M. Bieglmayer, Austria; N. Debry, Belgium; A. Dondelinger, Luxembourg; C. Ferrigno, Italy; T. R. Greve, Denmark; N. E. Haugen, Norway; T. Honohan, Ireland; J. V. Kahanpää, Finland; L. H. Kourtis, Greece; N. Leterrier, France; S. Mario del Fresno, Spain; C. R. B. Miller, U.K.; T. Moska, Sweden; J. P. de Carvalho Saraiva, Portugal; S. Schuh, Germany; U. N. Shimron, The Netherlands; E. Valenzuela, Chile; and M. Zavadsky, Switzerland) reports photometric observations with the 0.9-m Dutch telescope (CCD, standard B and V, ±0.1 mag) and spectroscopic observations with the 3.5-m New Technology Telescope (EMMI, 500-750 nm, resolution 0.3 nm) of the supernovae mentioned on IAUC 5895. 1993af. On Nov. 25.4 UT, B = 16.7, V = 16.6, B − V = +0.1. A 20-min spectrum on Nov. 27.3 showed this to be a type Ia supernova, 3-4 months after maximum, the age indicating that it is ∼ 4.5 mag fainter than at maximum. 1993ag. On Nov. 25.4, B = 18.1, V = 17.9, B −V = +0.2. A 30-min spectral exposure on Nov. 27.4 showed this supernova to be of type Ia, ∼ 1 week after maximum. A preliminary measurement of the expansion velocity, as deduced from the minimum of Si II absorption, gives ∼ 11 000 km/s. The redshift of the host galaxy, as measured from the Mg I absorption at 517 nm, is ∼ 0.07. 1993ah. On Nov. 25.1, B = 17.2, V = 16.4. B − V = +0.8. This supernova is situated in the outskirts of the central dustlane in the Sombrero-type galaxy ESO 471-G27. A 20-min spectrum on Nov. 27.1 showed it to be of type Ia, similar to SN 1981B ∼ 2-3 weeks after maximum. The expansion velocity, derived from the minimum of the Si II line, is ∼ 13 000 km/s.

PERIODIC COMET SPITALER (1993r)

Continuation to the ephemeris on IAUC 5885:

1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 29 1^{h}23.^{m}96 +13^{o}30.^{′}5 1.351 2.182 137.^{o}4 17.^{o}8 17.0
Dec. 9 1 24.74 +13 50.9 1.421 2.167 127.6 21.1 17.1
19 1 28.71 +14 24.7 1.504 2.155 118.6 23.6 17.2
29 1 35.69 +15 11.5 1.597 2.146 110.3 25.5 17.3
Jan. 8 1 45.34 +16 09.3 1.696 2.139 102.6 26.7 17.4
18 1 57.36 +17 16.1 1.801 2.135  95.6 27.3 17.6
28 2 11.44 +18 29.1 1.910 2.133  89.0 27.5 17.7

1993 November 27 Brian G. Marsden